OpenStudy (anonymous):

a worker uses a block and tackle with a mechanical advantage of 1.74 to lift a 225 N crate a distance of 16.5 m. He pulls 33.0 m of rope. What force did she use to lift the crate? 29.0 N , 129 N , 57.0 N , or 392 N

OpenStudy (michele_laino):

please, wait a moment a student asked me for help in Mathematics

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

here we have to consider the work done by the weight force of the crate, which is: \[{W_1} = Fs = 225 \times 16.5 = ...\]

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

i get 3712.5

OpenStudy (anonymous):

is that right?

OpenStudy (michele_laino):

ok! that's right!

OpenStudy (anonymous):

what happens next? to find the solution to the problem?

OpenStudy (michele_laino):

next we have to consider the work W_2 done by the worker, which is: \[{W_2} = Rd\] Now we don't know what is R

OpenStudy (michele_laino):

whereas d=330 meters

OpenStudy (michele_laino):

oops..d=33 meters

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

in order to find the value of the requasted force R, we have to apply this condition: \[\Large 1.74 = \frac{{3712.5}}{{33R}}\] so what is R?

OpenStudy (anonymous):

64.655?

OpenStudy (michele_laino):

sorry, it is the wrong result, please wait, I revise my reasoning

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

the requested force is given by the subsequent simple computation: \[1.74 = \frac{R}{F},\quad R = 1.74F = 1.74 \times 225 = ...Newtons\] here I have applied the definition, of advantage of a simple machine, namely: \[{\text{advantage}} = \frac{{{\text{resistive force}}}}{{{\text{agent force}}}}\]

OpenStudy (anonymous):

391.5? so our solution is 392 N?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yay!

OpenStudy (michele_laino):

:)